### v1

# A DRL-Empowered Multi-Level Jamming Approach for Secure Semantic Communication

Abstract
Semantic communication (SemCom) aims to transmit only task-relevant information, thereby improving communication efficiency but also exposing semantic information to potential eavesdropping. In this paper, we propose a deep reinforcement learning (DRL)-empowered multi-level jamming approach to enhance the security of SemCom systems over MIMO fading wiretap channels. This approach combines semantic layer jamming, achieved by encoding task-irrelevant text, and physical layer jamming, achieved by encoding random Gaussian noise. These two-level jamming signals are superposed with task-relevant semantic information to protect the transmitted semantics from eavesdropping. A deep deterministic policy gradient (DDPG) algorithm is further introduced to dynamically design and optimize the precoding matrices for both task-relevant semantic information and multi-level jamming signals, aiming to enhance the legitimate user’s image reconstruction while degrading the eavesdropper’s performance. To jointly train the SemCom model and the DDPG agent, we propose an alternating optimization strategy where the two modules are updated iteratively. Experimental results demonstrate that, compared with both the encryption-based (ESCS) and encoded jammer-based (EJ) benchmarks, our method achieves comparable security while improving the legitimate user’s peak signal-to-noise ratio (PSNR) by up to approximately 0.6 dB.
Index Terms:
Semantic communication, deep reinforcement learning, multi-level jamming, MIMO fading wiretap channels.
I Introduction
Semantic communication (SemCom) has emerged as a promising paradigm that improves communication efficiency by transmitting only task-relevant information, rather than raw symbols. Significant progress has been made in various data modalities, including text [1], image [2], speech [3], and video [4]. However, by focusing primarily on the essential semantics, SemCom typically eliminates traditional channel coding, which adds redundancy to improve reliability and provide implicit security. The absence of this redundancy introduces a vulnerability: transmitted data becomes more susceptible to eavesdropping, as intercepted signals can directly reveal sensitive content. As a result, secure SemCom over wiretap channels has become a crucial research direction and has recently attracted considerable attention [5]. Existing approaches can be broadly categorized into three categories: adversarial training-based methods, encryption-based methods, and physical layer-based methods.
From the adversarial training perspective, Erdemir et al. [6] proposed a VAE-based joint source-channel coding (JSCC) scheme that enhances security by maximizing the mutual information (MI) between the source and the legitimate channel output while penalizing information leakage. However, its security relies on variational approximations of MI, which may reduce its robustness against stronger adversaries. From the key-sharing-based encryption perspective, Luo et al. [7] developed an encrypted SemCom system (ESCS) that integrates a semantics-based symmetric cryptosystem learned through neural networks (NNs), and further proposed an adversarial encryption training strategy. However, it requires a secure channel for key exchange, which remains a limitation for practical deployment. More recent efforts have explored the use of homomorphic encryption in SemCom. Meng et al. [8] showed that semantic features can be retained in homomorphically encrypted data and proposed a privacy-preserved DeepJSCC framework, where activation and pooling functions are modified to support homomorphic operations.
From the physical layer perspective, recent efforts adopt Shannon capacity and secrecy capacity as fundamental design principles. Li et al. [9] proposed a two-phase training framework that balances reliable semantic recovery and reduced information leakage. The training process employs loss functions derived from Shannon and secrecy capacities, while a secure BLEU metric is used to evaluate semantic security against eavesdropping. Beyond capacity-based designs, other works have explored the use of artificial noise to enhance security. He et al. [10] proposed a framework where artificial noise is added at the transmitter and removed using denoising diffusion models at the legitimate receiver, thereby achieving a trade-off among communication reliability, privacy protection, and covertness. However, the use of generative denoising introduces considerable computational complexity and latency. Considering explicitly controllable system security, Chen et al. [11] superposed semantic information and random data on layered constellations under a tunable power allocation coefficient (PAC), thereby ensuring reliable decoding for the legitimate user and near-random decoding for the eavesdropper. However, the security of this method depends on a sufficient channel signal-to-noise ratio (SNR) gap between the legitimate user and the eavesdropper.
To address the limitations of existing approaches, we propose a novel deep reinforcement learning (DRL) [12]-empowered multi-level jamming framework for secure SemCom over MIMO fading wiretap channels. The main contributions of this paper are as follows.
  • We propose a novel secure SemCom framework that innovatively introduces a multi-level jamming mechanism. Unlike existing methods that typically provide protection at only a single level, our framework simultaneously incorporates semantic layer and physical layer jamming signals to protect the semantic information, offering more comprehensive and robust privacy protection.
  • We pioneer the integration of DRL into secure SemCom systems, which, to the best of our knowledge, has not been thoroughly explored in existing studies. Considering the complexity and challenges of the channel setting, we employ DRL to optimize the design of precoding matrices. These matrices empower the effective superposition and transmission of semantic information and jamming signals, thereby enhancing both task performance and system security.
  • Our framework addresses a highly challenging MIMO fading wiretap channel scenario, where both users experience comparable channel conditions, and Eve is assumed to have access to Bob’s receiver architecture and training dataset. Furthermore, unlike existing methods, our approach does not rely on secret key exchange, shared prior knowledge, or channel advantage, thereby reducing communication overhead and enhancing practicality. Moreover, the DRL agent introduces no additional computational overhead during inference.
  • We conduct extensive experiments to demonstrate that, compared with the key-sharing-based encrypted SemCom benchmark (ESCS) [7] and the encoded jammer-based benchmark (EJ) [13], our approach achieves comparable security levels while improving the peak signal-to-noise ratio (PSNR) performance of the legitimate user by up to approximately 0.6 dB.
II System Model
Figure 1: The framework of our proposed secure SemCom system.
In this paper, we consider a SemCom system designed for secure image transmission over MIMO fading wiretap channels, as illustrated in Fig. 1. We consider a MIMO SemCom system with $N_{m}\times N_{n}$ antennas. The system involves three entities: a transmitter (Alice), a legitimate receiver (Bob), and an eavesdropper (Eve). Alice aims to transmit a private source image X to Bob over MIMO fading channels so that Bob can accurately reconstruct the image while effectively preventing any unauthorized users from recovering it. Eve passively intercepts the semantic information through comparable MIMO fading channels and attempts to reconstruct X.
In addition to the source image $\textbf{X}\in\mathbb{R}^{H\times W\times C}$, other data is also included as input to the transmitter. We denote the entire input as Z. Alice employs a transmitter to encode the input Z in order to protect X, generating the semantic information to be transmitted, denoted as $\textbf{Y}\in\mathbb{R}^{N_{m}\times L_{c}}$:
$\textbf{Y}=f_{\mathrm{Alice}}\left(\textbf{Z}\right),$ (1)
where $f_{\mathrm{Alice}}$ represents the transmitter.
Subsequently, Y is normalized to satisfy the average transmit power constraint $P$, resulting in the normalized signal $\tilde{\textbf{Y}}$. This signal $\tilde{\textbf{Y}}$ is then transmitted through MIMO Rayleigh fading channels to Bob. Meanwhile, Eve intercepts the transmitted semantic information through her own channels. In this paper, we assume that the channel conditions for both the legitimate user and the eavesdropper are comparable. Accordingly, the semantic information received by Bob and Eve are denoted as ${\textbf{Y}}_{1}\in\mathbb{R}^{N_{n}\times L_{c}}$ and ${\textbf{Y}}_{2}\in\mathbb{R}^{N_{n}\times L_{c}}$, respectively, and are expressed as
${\textbf{Y}}_{1/2}=\textbf{H}\tilde{\textbf{Y}}+\textbf{N}_{1/2},$ (2)
where $\textbf{H}\in\mathbb{R}^{N_{n}\times N_{m}}$ denotes the channel matrix between Alice and Bob/Eve. Considering a conventional setting, we assume $N_{m}=N_{n}$. The channel coefficients h follow the distribution $\textbf{h}\sim\mathcal{N}(0,\textbf{I}_{N_{n}})$, where $\textbf{I}_{N_{n}}$ denotes the $N_{n}\times N_{n}$ identity matrix. $\textbf{N}_{1/2}\in\mathbb{R}^{N_{n}\times L_{c}}$ represents the additive white Gaussian noise (AWGN) term for Bob/Eve, where each element follows a Gaussian distribution $\mathcal{N}(0,\sigma_{1/2}^{2})$. The channel SNR between Alice and Bob/Eve can be calculated by $\mathrm{SNR}_{\rm leg/\rm eve}=10\log_{10}\left(\frac{P}{\sigma_{1/2}^{2}}\right)\,(\mathrm{dB})$.
At the receiver, before feeding the received semantic information ${\textbf{Y}}_{1}$ and ${\textbf{Y}}_{2}$ into their respective semantic decoders, Bob and Eve first perform signal detection on the received signals using minimum mean-squared error (MMSE) equalization [14]. MMSE equalization operates based on the estimated CSI $\hat{\textbf{H}}\in\mathbb{R}^{N_{n}\times N_{m}}$. In this paper, the estimated CSI is assumed to be perfect. The MMSE equalization process is as follows:
$\hat{\textbf{Y}}_{1/2}={\hat{\textbf{H}}}^{H}\left(\hat{\textbf{H}}{\hat{\textbf{H}}}^{H}+\frac{\sigma_{1/2}^{2}}{P}\textbf{I}\right)^{-1}\textbf{Y}_{1/2}.$ (3)
Here, $\hat{\textbf{Y}}_{1}$ and $\hat{\textbf{Y}}_{2}$ denote the MMSE-equalized versions of ${\textbf{Y}}_{1}$ and ${\textbf{Y}}_{2}$, respectively. $\sigma_{1/2}^{2}$ denotes the variance of the AWGN, and $P$ denotes the transmit power constraint.
After that, Bob uses his own receiver to reconstruct the source image based on $\hat{\textbf{Y}}_{1}$, resulting in $\hat{\textbf{X}}_{1}$:
$\hat{\textbf{X}}_{1}=f_{\mathrm{Bob}}\left(\hat{\textbf{Y}}_{1}\right),$ (4)
where $f_{\mathrm{Bob}}$ denotes Bob’s receiver, and $\hat{\textbf{X}}_{1}$ denotes the recovered image of Bob. Similarly, Eve uses her own receiver to reconstruct the source image based on $\hat{\textbf{Y}}_{2}$, resulting in $\hat{\textbf{X}}_{2}$:
$\hat{\textbf{X}}_{2}=f_{\mathrm{Eve}}\left(\hat{\textbf{Y}}_{2}\right),$ (5)
where $f_{\mathrm{Eve}}$ denotes Eve’s receiver, and $\hat{\textbf{X}}_{2}$ denotes the recovered image of Eve.
III Proposed Method
III-A Transmitter
The input Z to the transmitter consists of three components: the source image X, the text data T, and a randomly generated Gaussian noise matrix $\textbf{G}\in\mathbb{R}^{H\times W\times C}$. Each element in G follows a standard Gaussian distribution $\mathcal{N}(0,1)$. To protect the transmission of the source image X, a multi-level jamming scheme is employed at the transmitter. Specifically, the text data T, which is semantically irrelevant to X, is encoded to serve as the jamming signal at the semantic layer. Meanwhile, the Gaussian noise G is encoded to serve as the jamming signal at the physical layer. These two different levels of jamming can effectively protect the transmission of X and help reduce the risk of information leakage.
Specifically, at the transmitter, Alice first encodes the source image X to generate its semantic information $\textbf{S}_{1}$, which is formulated as
$\textbf{S}_{1}=f_{\mathrm{SE}}\left(\textbf{X};\bm{\theta}^{\mathrm{SE}}\right),$ (6)
where $f_{\mathrm{SE}}$ denotes the semantic encoder, and $\bm{\theta}^{\mathrm{SE}}$ represents its learnable parameters.
Subsequently, Alice encodes the text data T using a text jamming encoder (TJE) to generate the jamming signal at the semantic layer, denoted by $\textbf{S}_{2}$, which is formulated as
$\textbf{S}_{2}=f_{\mathrm{TJE}}\left(\textbf{T};\bm{\theta}^{\mathrm{TJE}}\right),$ (7)
where $f_{\mathrm{TJE}}$ denotes the text jamming encoder, and $\bm{\theta}^{\mathrm{TJE}}$ represents its learnable parameters.
Finally, Alice encodes the Gaussian noise G using a Gaussian jamming encoder (GJE) to generate the jamming signal at the physical layer, denoted by $\textbf{S}_{3}$, which is formulated as
$\textbf{S}_{3}=f_{\mathrm{GJE}}\left(\textbf{G};\bm{\theta}^{\mathrm{GJE}}\right),$ (8)
where $f_{\mathrm{GJE}}$ denotes the Gaussian jamming encoder, and $\bm{\theta}^{\mathrm{GJE}}$ represents its learnable parameters.
After generating the semantic information and multi-level jamming signals, Alice applies the precoding matrices $\textbf{V}_{1}$, $\textbf{V}_{2}$, and $\textbf{V}_{3}$, which are produced by the DDPG agent (to be discussed later), to precode the signals $\textbf{S}_{1}$, $\textbf{S}_{2}$, and $\textbf{S}_{3}$, respectively. These precoded signals are then combined through superposition coding to form the transmitted semantic information, denoted as Y. The process is described by the following equation:
$\textbf{Y}=\textbf{V}_{1}\textbf{S}_{1}+\textbf{V}_{2}\textbf{S}_{2}+\textbf{V}_{3}\textbf{S}_{3},$ (9)
where $\textbf{V}_{1}$, $\textbf{V}_{2}$, and $\textbf{V}_{3}$ are each of dimension $N_{m}\times N_{n}$, and $\textbf{S}_{1}$, $\textbf{S}_{2}$, and $\textbf{S}_{3}$ are each of dimension $N_{n}\times L_{c}$. As a result, Y has a dimension of $N_{m}\times L_{c}$.
III-B Receivers at Bob and Eve
In this paper, we assume that Eve has obtained the network architecture of Bob’s semantic decoder. For Bob, the MMSE-equalized received signal $\hat{\textbf{Y}}_{1}$ is input into his semantic decoder to reconstruct the image $\hat{\textbf{X}}_{1}$:
$\hat{\textbf{X}}_{1}=f_{\mathrm{SD1}}\left(\hat{\textbf{Y}}_{1};\bm{\theta}^{\mathrm{SD1}}\right),$ (10)
where $f_{\mathrm{SD1}}$ denotes Bob’s semantic decoder, and $\bm{\theta}^{\mathrm{SD1}}$ represents its learnable parameters.
Similarly, Eve inputs the MMSE-equalized received signal $\hat{\textbf{Y}}_{2}$ into her semantic decoder to reconstruct the image $\hat{\textbf{X}}_{2}$:
$\hat{\textbf{X}}_{2}=f_{\mathrm{SD2}}\left(\hat{\textbf{Y}}_{2};\bm{\theta}^{\mathrm{SD2}}\right),$ (11)
where $f_{\mathrm{SD2}}$ denotes Eve’s semantic decoder, and $\bm{\theta}^{\mathrm{SD2}}$ represents its learnable parameters.
III-C The DDPG Agent
We consider a standard DRL setup that includes a DDPG agent, which observes the environment in discrete time slots. During the training stage of the proposed system, the DDPG agent interacts with the environment to obtain the current state, which is then fed into the action network to generate three precoding matrices for the transmitter. After the environment executes the action policy and undergoes training for a certain number of epochs, the DDPG agent receives feedback from the updated environment, evaluates the value of its action policy, and optimizes the parameters of its networks. As shown in Fig. 1, the DDPG agent consists of an actor network and a critic network. In addition, there are target networks, namely the target actor network and the target critic network, which serve as copies of the actor and critic networks, respectively. These target networks are softly updated by getting parameters from their corresponding main networks.
We begin by introducing the state-action definitions for the proposed secure SemCom system.
  • State: At each decision time step $t$, the environment state $\textbf{s}_{t}\in\mathbb{R}^{7}$ consists of the following variables:
    $\begin{split}\textbf{s}_{t}=\left[{\rm SNR}^{\rm leg}_{t},{\rm SNR}^{\rm eve}_{t},CU,\right.\\ \left.{\rm PSNR}^{\rm leg}_{t-1},{\rm PSNR}^{\rm eve}_{t-1},\mathcal{L}_{t-1},t/T\right],\end{split}$ (12)
    where ${\rm SNR}^{\rm leg}_{t}$ and ${\rm SNR}^{\rm eve}_{t}$ denote the SNRs of the legitimate user and the eavesdropper, respectively. $CU$ controls the compression ratio of the system. ${\rm PSNR}^{\rm leg}_{t-1}$ and ${\rm PSNR}^{\rm eve}_{t-1}$ indicate the image reconstruction quality (PSNR) from the previous time step for Bob and Eve, respectively. $\mathcal{L}_{t-1}$ denotes the previous loss, and $t/T$ represents the relative training progress.
  • Action: At each decision time step $t$, the action $\textbf{a}_{t}\in\mathbb{R}^{3\cdot N_{m}\cdot N_{n}}$ is defined as the output of the actor network. It represents the flattened form of three precoding matrices $\{\textbf{V}_{1},\textbf{V}_{2},\textbf{V}_{3}\}$, each of dimension $N_{m}\times N_{n}$. The action is subsequently reshaped into a tensor of size $3\times N_{m}\times N_{n}$ as:
    $\textbf{a}_{t}\Rightarrow\left[\textbf{V}_{1},\textbf{V}_{2},\textbf{V}_{3}\right],\textbf{V}_{i}\in\mathbb{R}^{N_{m}\times N_{n}},\;i=1,2,3.$ (13)
  • Reward: The reward $r_{t}\in\mathbb{R}$ at each decision time step is designed to simultaneously encourage high image quality for the legitimate user while limiting information leakage to the eavesdropper. It is defined as:
    $r_{t}=\mathrm{PSNR}^{\mathrm{leg}}_{t}-\lambda_{r}\cdot\mathrm{PSNR}^{\mathrm{eve}}_{t},$ (14)
    where $\mathrm{PSNR}^{\mathrm{leg}}_{t}$ and $\mathrm{PSNR}^{\mathrm{eve}}_{t}$ denote the image reconstruction quality for Bob and Eve, respectively. $\lambda_{r}$ is a tunable parameter that balances the trade-off between task performance and system security.
Subsequently, we describe how the DDPG agent is optimized to design effective precoding strategies. We define the actor network $\mu_{\theta}$ and the critic network $Q_{\phi}$. Both networks are composed of several fully connected (linear) layers. To stabilize the training process, target networks $\mu_{\theta^{\prime}}$ and $Q_{\phi^{\prime}}$ are maintained as time-delayed copies of the actor and critic networks, respectively. A replay buffer $\mathcal{D}$ is used to store past transition tuples $(\textbf{s}_{t},\textbf{a}_{t},r_{t},\textbf{s}_{t+1},d_{t})$, where $d_{t}$ denotes a terminal signal. At each decision time step, exploration noise sampled from an Ornstein-Uhlenbeck (OU) process $\mathcal{N}_{t}$ is added to the actor’s deterministic output to encourage temporally correlated exploration in continuous action space:
$\textbf{a}_{t}=\mu_{\theta}(\textbf{s}_{t})+\mathcal{N}_{t}.$ (15)
The critic network is optimized by minimizing the mean squared Bellman error over a minibatch of $B$ experiences sampled from $\mathcal{D}$:
$\mathcal{L}_{\text{critic}}=\frac{1}{B}\sum_{i=1}^{B}\left(y_{i}-Q_{\phi}(\textbf{s}_{i},\textbf{a}_{i})\right)^{2},$ (16)
where the temporal-difference target $y_{i}$ is computed using the target actor and critic networks as:
$y_{i}=r_{i}+\gamma\cdot Q_{\phi^{\prime}}(\textbf{s}_{i+1},\mu_{\theta^{\prime}}(\textbf{s}_{i+1}))\cdot(1-d_{i}),$ (17)
with $\gamma$ denoting the reward discount factor.
The actor network is updated by applying the deterministic policy gradient, which aims to maximize the expected critic value under the current policy:
$\nabla_{\theta}\mathcal{J}\approx\frac{1}{B}\sum_{i=1}^{B}\nabla_{\theta}\mu_{\theta}(\textbf{s}_{i})\nabla_{\tilde{\textbf{a}}}Q_{\phi}(\textbf{s}_{i},\tilde{\textbf{a}})\big|_{\tilde{\textbf{a}}=\mu_{\theta}(\textbf{s}_{i})}.$ (18)
Accordingly, the following loss function is minimized in practice:
$\mathcal{L}_{\text{actor}}=-\frac{1}{B}\sum_{i=1}^{B}Q_{\phi}(\textbf{s}_{i},\mu_{\theta}(\textbf{s}_{i})).$ (19)
To ensure stable learning, the parameters of the target networks are updated via soft updates after each gradient step.
III-D Training Strategy
In this paper, we propose a five-stage training strategy to train the proposed system.
III-D1 Stage 1
In this training stage, we focus on training the semantic encoder $f_{\text{SE}}$ and the semantic decoder of the legitimate user $f_{\text{SD1}}$, while keeping all other modules frozen. The transmitted semantic information in this stage is given by $\textbf{Y}=\textbf{S}_{1}$. The system is optimized using
$\mathcal{L}_{1}={\rm MSE}(\textbf{X},\hat{\textbf{X}}_{1}).$ (20)
III-D2 Stage 2
In this training stage, multi-level jamming is introduced into the system, but the DDPG agent is not yet involved. All modules are activated except for the eavesdropper’s decoder and the DDPG agent. The transmitted semantic information in this stage is given by $\textbf{Y}=\textbf{S}_{1}+\textbf{S}_{2}+\textbf{S}_{3}$. The loss function remains:
$\mathcal{L}_{2}={\rm MSE}(\textbf{X},\hat{\textbf{X}}_{1}).$ (21)
III-D3 Stage 3
In this training stage, the DDPG agent remains inactive. The eavesdropper’s decoder $f_{\text{SD2}}$ is trained independently, with all other modules kept frozen. The transmitted semantic information in this stage is given by $\textbf{Y}=\textbf{S}_{1}+\textbf{S}_{2}+\textbf{S}_{3}$. The loss function used is:
$\mathcal{L}_{3}={\rm MSE}(\textbf{X},\hat{\textbf{X}}_{2}),$ (22)
where $\hat{\textbf{X}}_{2}$ denotes Eve’s reconstructed image.
III-D4 Stage 4
In this training stage, the DDPG agent is introduced and jointly trained with the SemCom model (environment) through an alternating update scheme. The transmitted semantic information in this stage is given by $\textbf{Y}=\textbf{V}_{1}\textbf{S}_{1}+\textbf{V}_{2}\textbf{S}_{2}+\textbf{V}_{3}\textbf{S}_{3}$. Specifically, the training alternates between two stages:
  • Agent update stage: At the beginning of every $K$ epochs, the DDPG agent observes the current environment state and generates a new action $\textbf{a}_{t}$, corresponding to a set of precoding matrices $\{\textbf{V}_{1},\textbf{V}_{2},\textbf{V}_{3}\}$. This action remains fixed throughout the subsequent $K$ epochs of environment training. After receiving the reward and state transition at the end of the $K$ epochs, the agent stores the resulting experience tuple in its replay buffer. If the buffer contains enough samples, the agent updates its actor and critic networks via mini-batch gradient descent.
  • Environment update stage: For the next $K$ epochs, the SemCom model is trained using the fixed action provided by the agent. During this stage, all modules are updated except for the eavesdropper’s decoder $f_{{\rm SD2}}$. After $K$ epochs, the environment reports the reward and updated state to the agent, which then uses this feedback to refine its policy.
This alternating training procedure continues until the predefined number of training epochs is completed. The loss function used is:
$\mathcal{L}_{4}={\rm MSE}(\textbf{X},\hat{\textbf{X}}_{1})-\lambda_{r}\cdot{\rm MSE}(\textbf{X},\hat{\textbf{X}}_{2}),$ (23)
where $\lambda_{r}$ is the same as in equation (14).
III-D5 Stage 5
In this training stage, the optimal policy obtained in stage 4, selected based on the highest reward, is saved and directly applied to guide both the training and testing of the final system. The transmitted semantic information is generated in the same way as in stage 4, and the same loss function is used.
IV Performance Evaluation
IV-A Experimental Settings
IV-A1 Datasets
We use the CIFAR-10 dataset for the source image X. For the text data T, we utilize the proceedings of the European Parliament dataset [15].
IV-A2 Training Settings
We set $N_{m}=N_{n}=4$. Both users are assumed to experience the same channel SNR, given by $\mathrm{SNR}_{\mathrm{leg}}=\mathrm{SNR}_{\mathrm{eve}}\in\{0,5,10,15,20\}\,\mathrm{dB}$. The value of compression ratio (CR) is controlled by the variable $CU$. When $CU=1$, the corresponding CR is $1/96$. The CR takes values from the set $CR\in\{1/96,2/96,3/96,4/96,5/96\}$.
For the DDPG agent, $K=15$, the total number of decision time steps $T$ is set to 500, and the trade-off coefficient $\lambda_{r}$ is set to 1.0. The replay buffer size is set to 1000, and the minibatch size $B$ is 128. The discount factor $\gamma$ is set to 0.99, and the soft update coefficient $\tau$ is set to $1\times 10^{-3}$. The learning rate of the actor network is $3\times 10^{-4}$, and that of the critic network is also $3\times 10^{-4}$. The L2 weight decay is set to $1\times 10^{-4}$.
The learning rate in the first three training stages is set to $1\times 10^{-3}$ and maintained for 100 epochs. In the fourth training stage, the learning rate is set to $5\times 10^{-4}$ and maintained for $T\cdot K$ epochs, which correspond to 7500 epochs. In the fifth training stage, the learning rate is set to $2\times 10^{-4}$ and maintained for 200 epochs.
IV-A3 The Benchmarks
The first benchmark is the ESCS scheme [7], which is a key-based encryption approach designed for secure SemCom systems over wiretap channels. The second benchmark is the EJ scheme [13], which is a jamming-based scheme developed for secure communication systems over MIMO wiretap channels. Both benchmarks employ the classical singular value decomposition (SVD) precoding and MMSE equalization.
IV-B Performance Comparison of Different Approaches
IV-B1 Different Channel SNRs
Figure 2: Comparison of PSNR performance for both users under different channel SNRs using different approaches. The compression ratio is fixed at 1/96, with $N_{m}$ and $N_{n}$ both set to 4. We set $\rm SNR_{leg}=\rm SNR_{eve}$.
In this subsection, we evaluate the PSNR performance of the proposed system under varying channel SNRs. The experimental results are shown in Fig. 2, with a fixed CR of 1/96. As shown in Fig. 2, the PSNR of the legitimate user steadily improves as the channel SNR increases. Our proposed method consistently outperforms both EJ and ESCS across all channel SNRs, providing a notable performance gain due to its efficient DRL-empowered multi-level jamming strategy. Meanwhile, the eavesdropper’s PSNR remains at a consistently low level and exhibits only minor fluctuations with changes in channel SNR, indicating that the superposed semantic and physical layer jamming effectively prevents accurate image reconstruction for the eavesdropper. While ESCS achieves slightly lower PSNR for the eavesdropper, our proposed method attains a more favorable trade-off by enhancing the legitimate user’s reconstruction quality without significantly compromising security.
Figure 3: Comparison of PSNR performance for both users under different compression ratios using different approaches. We set $\rm SNR_{leg}=\rm SNR_{eve}=10dB$, with $N_{m}$ and $N_{n}$ both set to 4.
IV-B2 Different Compression Ratios
In this subsection, we evaluate the PSNR performance of the proposed system under different compression ratios and compare it with the two benchmarks. As shown in Fig. 3, the PSNR of the legitimate user increases steadily as the compression ratio grows. Among all evaluated methods, the proposed method consistently achieves the highest reconstruction quality across all compression ratios. Specifically, at a compression ratio of 5/96, our method attains a PSNR of 20.92 dB, outperforming EJ (20.14 dB) and ESCS (20.54 dB). In contrast to the legitimate user, the eavesdropper’s PSNR exhibits a decreasing trend with higher compression ratios. At the highest compression ratio, our method reduces the eavesdropper’s PSNR to 6.03 dB, which is lower than that of EJ (6.46 dB) and ESCS (6.20 dB), indicating improved system security. Overall, compared with the benchmarks, the proposed system consistently achieves a superior trade-off between task performance and system security, demonstrating high effectiveness and robustness under varying compression ratios.
V Conclusion
In this paper, we proposed a SemCom framework that integrated a DRL-empowered multi-level jamming strategy to enhance transmission security over MIMO fading wiretap channels. The framework was designed for a realistic and challenging scenario where both the legitimate user and the eavesdropper experienced comparable channel conditions, without relying on secret key exchange or any prior shared knowledge. Specifically, two levels of jamming signals were introduced: semantic layer jamming, generated by encoding task-irrelevant text data, and physical layer jamming, generated by encoding random Gaussian noise. These jamming signals were superposed with task-relevant semantic information to effectively obscure the transmitted semantics from the eavesdropper while maintaining high reconstruction quality for the legitimate user. To optimize and empower this process, we employed the DDPG algorithm to design and adapt the precoding matrices for both task-relevant semantic information and multi-level jamming signals. This approach effectively balanced the trade-off between the legitimate user’s reconstruction quality and overall system security. Moreover, the DDPG agent operated only during training, introducing no additional computational overhead during inference. Experimental results demonstrated that, compared with the ESCS and EJ benchmarks, our approach achieved comparable security while improving the legitimate user’s PSNR by up to approximately 0.6 dB. These results confirmed that the proposed DRL-empowered multi-level jamming framework provided an effective and practical solution for secure SemCom systems operating under challenging conditions.

### v3

@@
 As a result, secure SemCom over wiretap channels has become a crucial research direction and has recently attracted considerable attention [4].
-Existing approaches can be broadly categorized into three categories: adversarial training-based methods, encryption-based methods, and physical layer-based methods.
+Existing approaches can be broadly categorized into three groups: adversarial training-based methods, encryption-based methods, and physical layer-based methods.
 From the adversarial training perspective, Erdemir et al.
-[6] proposed a VAE-based joint source-channel coding (JSCC) scheme that enhances security by maximizing the mutual information (MI) between the source and the legitimate channel output while penalizing information leakage.
+[6] proposed a variational autoencoder (VAE)-based deep joint source-channel coding (DeepJSCC) scheme that enhances security by maximizing the mutual information (MI) between the source and the legitimate channel output while penalizing information leakage.
 However, its security relies on variational approximations of MI, which may reduce its robustness against stronger adversaries.
@@
 Eve passively intercepts the semantic information through comparable MIMO fading channels and attempts to reconstruct X.
-In addition to the source image $\textbf{X}\in\mathbb{R}^{H\times W\times C}$, other data is also included as input to the transmitter.
+In addition to the source image $\textbf{X}\in\mathbb{R}^{H\times W\times C}$, other information is also included as input to the transmitter.
 We denote the entire input as Z.
@@
 where $f_{\mathrm{Alice}}$ represents the transmitter.
-Subsequently, Y is normalized to satisfy the average transmit power constraint $P$, resulting in the normalized signal $\tilde{\textbf{Y}}$.
+Subsequently, Y is normalized with respect to its average power in order to satisfy the transmit power constraint $P$ and then converted into a complex-valued representation, resulting in the normalized signal $\tilde{\textbf{Y}}\in\mathbb{C}^{N_{m}\times\frac{L_{c}}{2}}$.
 This signal $\tilde{\textbf{Y}}$ is then transmitted through MIMO Rayleigh fading channels to Bob.
 Meanwhile, Eve intercepts the transmitted semantic information through her own channels.
-In this paper, we assume that the channel conditions for both the legitimate user and the eavesdropper are comparable.
-Accordingly, the semantic information received by Bob and Eve are denoted as ${\textbf{Y}}_{1}\in\mathbb{R}^{N_{n}\times L_{c}}$ and ${\textbf{Y}}_{2}\in\mathbb{R}^{N_{n}\times L_{c}}$, respectively, and are expressed as
+Accordingly, the semantic information received by Bob and Eve are denoted as ${\textbf{Y}}_{1}\in\mathbb{C}^{N_{n}\times\frac{L_{c}}{2}}$ and ${\textbf{Y}}_{2}\in\mathbb{C}^{N_{n}\times\frac{L_{c}}{2}}$, respectively, and are expressed as
 ${\textbf{Y}}_{1/2}=\textbf{H}\tilde{\textbf{Y}}+\textbf{N}_{1/2},$ (2)
-where $\textbf{H}\in\mathbb{R}^{N_{n}\times N_{m}}$ denotes the channel matrix between Alice and Bob/Eve.
+where $\textbf{H}\in\mathbb{C}^{N_{n}\times N_{m}}$ denotes the channel matrix between Alice and Bob/Eve.
 Considering a conventional setting, we assume $N_{m}=N_{n}$.
-The channel coefficients h follow the distribution $\textbf{h}\sim\mathcal{N}(0,\textbf{I}_{N_{n}})$, where $\textbf{I}_{N_{n}}$ denotes the $N_{n}\times N_{n}$ identity matrix.
-$\textbf{N}_{1/2}\in\mathbb{R}^{N_{n}\times L_{c}}$ represents the additive white Gaussian noise (AWGN) term for Bob/Eve, where each element follows a Gaussian distribution $\mathcal{N}(0,\sigma_{1/2}^{2})$.
+The channel coefficients h follow the distribution $\textbf{h}\sim\mathcal{CN}(0,\textbf{I}_{N_{n}})$, where $\textbf{I}_{N_{n}}$ denotes the $N_{n}\times N_{n}$ identity matrix.
+$\textbf{N}_{1/2}\in\mathbb{C}^{N_{n}\times\frac{L_{c}}{2}}$ represents the additive white Gaussian noise (AWGN) term for Bob/Eve, where each element follows a complex Gaussian distribution $\mathcal{CN}(0,\sigma_{1/2}^{2})$.
 The channel SNR between Alice and Bob/Eve can be calculated by $\mathrm{SNR}_{\rm leg/\rm eve}=10\log_{10}\left(\frac{P}{\sigma_{1/2}^{2}}\right)\,(\mathrm{dB})$.
 At the receiver, before feeding the received semantic information ${\textbf{Y}}_{1}$ and ${\textbf{Y}}_{2}$ into their respective semantic decoders, Bob and Eve first perform signal detection on the received signals using minimum mean-squared error (MMSE) equalization [15].
-MMSE equalization operates based on the estimated CSI $\hat{\textbf{H}}\in\mathbb{R}^{N_{n}\times N_{m}}$.
+MMSE equalization operates based on the estimated channel state information (CSI) $\hat{\textbf{H}}\in\mathbb{C}^{N_{n}\times N_{m}}$.
 In this paper, the estimated CSI is assumed to be perfect.
 The MMSE equalization process is as follows:
-$\hat{\textbf{Y}}_{1/2}={\hat{\textbf{H}}}^{H}\left(\hat{\textbf{H}}{\hat{\textbf{H}}}^{H}+\frac{\sigma_{1/2}^{2}}{P}\textbf{I}\right)^{-1}\textbf{Y}_{1/2}.$ (3)
-Here, $\hat{\textbf{Y}}_{1}$ and $\hat{\textbf{Y}}_{2}$ denote the MMSE-equalized versions of ${\textbf{Y}}_{1}$ and ${\textbf{Y}}_{2}$, respectively.
+$\bar{\textbf{Y}}_{1/2}={\hat{\textbf{H}}}^{H}\left(\hat{\textbf{H}}{\hat{\textbf{H}}}^{H}+\frac{\sigma_{1/2}^{2}}{P}\textbf{I}\right)^{-1}\textbf{Y}_{1/2}.$ (3)
+Here, $\bar{\textbf{Y}}_{1}$ and $\bar{\textbf{Y}}_{2}$ denote the MMSE-equalized versions of ${\textbf{Y}}_{1}$ and ${\textbf{Y}}_{2}$, respectively.
 $\sigma_{1/2}^{2}$ denotes the variance of the AWGN, and $P$ denotes the transmit power constraint.
+Subsequently, $\bar{\textbf{Y}}_{1}$ and $\bar{\textbf{Y}}_{2}$ are converted back to real-valued representations, denoted by $\hat{\textbf{Y}}_{1}$ and $\hat{\textbf{Y}}_{2}$.
 After that, Bob uses his own receiver to reconstruct the source image based on $\hat{\textbf{Y}}_{1}$, resulting in $\hat{\textbf{X}}_{1}$:
@@
 where $f_{\mathrm{GJE}}$ denotes the Gaussian jamming encoder, and $\bm{\theta}^{\mathrm{GJE}}$ represents its learnable parameters.
-After generating the semantic information and multi-level jamming signals, Alice applies the precoding matrices $\textbf{V}_{1}$, $\textbf{V}_{2}$, and $\textbf{V}_{3}$, which are produced by the DDPG agent (to be discussed later), to precode the signals $\textbf{S}_{1}$, $\textbf{S}_{2}$, and $\textbf{S}_{3}$, respectively.
+After generating the semantic information and multi-level jamming signals, Alice applies the precoding matrices $\textbf{V}_{1}$, $\textbf{V}_{2}$, and $\textbf{V}_{3}$, which are produced by the deep deterministic policy gradient (DDPG) agent (to be discussed later), to precode the signals $\textbf{S}_{1}$, $\textbf{S}_{2}$, and $\textbf{S}_{3}$, respectively.
 These precoded signals are then combined through superposition coding to form the transmitted semantic information, denoted as Y.
@@
   • State: At each decision time step $t$, the environment state $\textbf{s}_{t}\in\mathbb{R}^{7}$ consists of the following variables:
-    $\begin{split}\textbf{s}_{t}=\left[{\rm SNR}^{\rm leg}_{t},{\rm SNR}^{\rm eve}_{t},CU,\right.\\ \left.{\rm PSNR}^{\rm leg}_{t-1},{\rm PSNR}^{\rm eve}_{t-1},\mathcal{L}_{t-1},t/T\right],\end{split}$ (12)
+    $\begin{split}\textbf{s}_{t}=\left[{\rm SNR}^{\rm leg}_{t},{\rm SNR}^{\rm eve}_{t},CU,\right.\\ \left.{\rm PSNR}^{\rm leg}_{t-1},{\rm PSNR}^{\rm eve}_{t-1},\mathcal{L}_{t-1},\frac{t}{T}\right],\end{split}$ (12)
     where ${\rm SNR}^{\rm leg}_{t}$ and ${\rm SNR}^{\rm eve}_{t}$ denote the SNRs of the legitimate user and the eavesdropper, respectively.
-$CU$ controls the compression ratio of the system.
+$CU$ controls the compression ratio (CR) of the system.
 ${\rm PSNR}^{\rm leg}_{t-1}$ and ${\rm PSNR}^{\rm eve}_{t-1}$ indicate the image reconstruction quality (PSNR) from the previous time step for Bob and Eve, respectively.
-$\mathcal{L}_{t-1}$ denotes the previous loss, and $t/T$ represents the relative training progress.
+$\mathcal{L}_{t-1}$ denotes the previous loss, and $\frac{t}{T}$ represents the relative training progress.
   • Action: At each decision time step $t$, the action $\textbf{a}_{t}\in\mathbb{R}^{3\cdot N_{m}\cdot N_{n}}$ is defined as the output of the actor network.
@@
 III-D1 Stage 1
-In this training stage, we focus on training the semantic encoder $f_{\text{SE}}$ and the semantic decoder of the legitimate user $f_{\text{SD1}}$, while keeping all other modules frozen.
+In this stage, we focus on training the semantic encoder $f_{\text{SE}}$ and the semantic decoder of the legitimate user $f_{\text{SD1}}$, while keeping all other modules frozen.
 The transmitted semantic information in this stage is given by $\textbf{Y}=\textbf{S}_{1}$.
@@
 III-D2 Stage 2
-In this training stage, multi-level jamming is introduced into the system, but the DDPG agent is not yet involved.
+In this stage, multi-level jamming is introduced into the system, but the DDPG agent is not yet involved.
 All modules are activated except for the eavesdropper’s decoder and the DDPG agent.
@@
 III-D3 Stage 3
-In this training stage, the DDPG agent remains inactive.
+In this stage, the DDPG agent remains inactive.
 The eavesdropper’s decoder $f_{\text{SD2}}$ is trained independently, with all other modules kept frozen.
@@
 III-D4 Stage 4
-In this training stage, the DDPG agent is introduced and jointly trained with the SemCom model (environment) through an alternating update scheme.
+In this stage, the DDPG agent is introduced and jointly trained with the SemCom model (environment) through an alternating update scheme.
 The transmitted semantic information in this stage is given by $\textbf{Y}=\textbf{V}_{1}\textbf{S}_{1}+\textbf{V}_{2}\textbf{S}_{2}+\textbf{V}_{3}\textbf{S}_{3}$.
 Specifically, the training alternates between two stages:
-  • Agent update stage: At the beginning of every $K$ epochs, the DDPG agent observes the current environment state and generates a new action $\textbf{a}_{t}$, corresponding to a set of precoding matrices $\{\textbf{V}_{1},\textbf{V}_{2},\textbf{V}_{3}\}$.
+  • Agent update stage: At the beginning of each $K$-epoch interval, the DDPG agent observes the current environment state and generates a new action $\textbf{a}_{t}$, corresponding to a set of precoding matrices $\{\textbf{V}_{1},\textbf{V}_{2},\textbf{V}_{3}\}$.
 This action remains fixed throughout the subsequent $K$ epochs of environment training.
@@
 III-D5 Stage 5
-In this training stage, the optimal policy obtained in stage 4, selected based on the highest reward, is saved and directly applied to guide both the training and testing of the final system.
+In this stage, the optimal policy obtained in stage 4, selected based on the highest reward, is saved and directly applied to guide both the training and testing of the final system.
 The transmitted semantic information is generated in the same way as in stage 4, and the same loss function is used.
@@
 Both users are assumed to experience the same channel SNR, given by $\mathrm{SNR}_{\mathrm{leg}}=\mathrm{SNR}_{\mathrm{eve}}\in\{0,5,10,15,20\}\,\mathrm{dB}$.
-The value of compression ratio (CR) is controlled by the variable $CU$.
+The CR is controlled by the variable $CU$.
 When $CU=1$, the corresponding CR is $1/96$.
@@
 The learning rate in the first three training stages is set to $1\times 10^{-3}$ and maintained for 100 epochs.
-In the fourth training stage, the learning rate is set to $5\times 10^{-4}$ and maintained for $T\cdot K$ epochs, which correspond to 7500 epochs.
+In the fourth training stage, the learning rate is set to $5\times 10^{-4}$ and maintained for $T\cdot K$ epochs, which corresponds to 7500 epochs.
 In the fifth training stage, the learning rate is set to $2\times 10^{-4}$ and maintained for 200 epochs.
@@
 The second benchmark is the EJ scheme [14], which is a jamming-based scheme developed for secure communication systems over MIMO wiretap channels.
-Both benchmarks employ the classical singular value decomposition (SVD) precoding and MMSE equalization.
+In both benchmarks, we employ classical singular value decomposition (SVD) precoding and MMSE equalization.
 IV-B Performance Comparison of Different Approaches
@@
 Figure 2: Comparison of PSNR performance for both users under different channel SNRs using different approaches.
-The compression ratio is fixed at 1/96, with $N_{m}$ and $N_{n}$ both set to 4.
+The CR is fixed at $1/96$, with $N_{m}$ and $N_{n}$ both set to 4.
 We set $\rm SNR_{leg}=\rm SNR_{eve}$.
@@
 While ESCS achieves slightly lower PSNR for the eavesdropper, our proposed method attains a more favorable trade-off by enhancing the legitimate user’s reconstruction quality without significantly compromising security.
-Figure 3: Comparison of PSNR performance for both users under different compression ratios using different approaches.
-We set $\rm SNR_{leg}=\rm SNR_{eve}=10dB$, with $N_{m}$ and $N_{n}$ both set to 4.
+Figure 3: Comparison of PSNR performance for both users under different CRs using different approaches. We set $\rm SNR_{leg}=\rm SNR_{eve}=10dB$, with $N_{m}$ and $N_{n}$ both set to 4.
 IV-B2 Different Compression Ratios
-In this subsection, we evaluate the PSNR performance of the proposed system under different compression ratios and compare it with the two benchmarks.
+In this subsection, we evaluate the PSNR performance of the proposed system under different CRs and compare it with the two benchmarks.
 As shown in Fig.
-3, the PSNR of the legitimate user increases steadily as the compression ratio grows.
-Among all evaluated methods, the proposed method consistently achieves the highest reconstruction quality across all compression ratios.
-Specifically, at a compression ratio of 5/96, our method attains a PSNR of 20.92 dB, outperforming EJ (20.14 dB) and ESCS (20.54 dB).
-In contrast to the legitimate user, the eavesdropper’s PSNR exhibits a decreasing trend with higher compression ratios.
-At the highest compression ratio, our method reduces the eavesdropper’s PSNR to 6.03 dB, which is lower than that of EJ (6.46 dB) and ESCS (6.20 dB), indicating improved system security.
-Overall, compared with the benchmarks, the proposed system consistently achieves a superior trade-off between task performance and system security, demonstrating high effectiveness and robustness under varying compression ratios.
+3, the PSNR of the legitimate user increases steadily as the CR grows.
+Among all evaluated methods, the proposed method consistently achieves the highest reconstruction quality across all CRs.
+Specifically, at a CR of $5/96$, our method attains a PSNR of 20.92 dB, outperforming EJ (20.14 dB) and ESCS (20.54 dB).
+In contrast to the legitimate user, the eavesdropper’s PSNR exhibits a decreasing trend with higher CRs.
+At the highest CR, our method reduces the eavesdropper’s PSNR to 6.03 dB, which is lower than that of EJ (6.46 dB) and ESCS (6.20 dB), indicating improved system security.
+Overall, compared with the benchmarks, the proposed system consistently achieves a superior trade-off between task performance and system security, demonstrating high effectiveness and robustness under varying CRs.
 V Conclusion
@@
 Experimental results demonstrated that, compared with the ESCS and EJ benchmarks, our approach achieved comparable security while improving the legitimate user’s PSNR by up to approximately 0.6 dB.
-These results confirmed that the proposed DRL-empowered multi-level jamming framework provided an effective and practical solution for secure SemCom systems operating under challenging conditions.
+These results confirm that the proposed DRL-empowered multi-level jamming framework provides an effective and practical solution for secure SemCom systems operating under challenging conditions.